OpenStudy (anonymous):

Which statement is NOT true? A. If x = 10, then x2 = 100. B. x2 = 100 if and only if x = 10 or x = –10. C. If x = –10, then x2 = 100. D. If x2 = 100, then x = 10.

ganeshie8 (ganeshie8):

check the last one once

OpenStudy (anonymous):

D. Because, A. \[10 * 10 = 100\] B. \[10 * 10 = 100, -10 * - 10 = 100\] (negative times a negative.) C. Same as above, -10*-10 = 100 D. If x^2 = 100, then x can be -10, or 10.